OpenStudy (junyang96):

(a-b)^2=(b-a)^2 where a and b are real numbers is the statement ever correct?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

another way, a little wierd \(\large\color{black}{ \displaystyle (a-b)^2=(b-a)^2 }\) \(\large\color{black}{ \displaystyle \sqrt{ (a-b)^2}=\sqrt{(b-a)^2} }\) \(\large\color{black}{ \displaystyle |a-b|=|b-a| }\)

OpenStudy (solomonzelman):

or you know that let a - b = C and so b - a = -C you know (-C)^2=(C)^2

OpenStudy (junyang96):

cool thanks

OpenStudy (solomonzelman):

yes, the first approach is like a distance from "a" to "b" and the second is noticing that \(\large\color{black}{ \displaystyle x^2=a^2 }\) then \(\large\color{black}{ \displaystyle x=\pm a }\)

OpenStudy (solomonzelman):

you welcome, but remember this that the definition of the absolute value of f(x) is \(\large\color{black}{ \displaystyle |f(x)|=\sqrt{ \left[f(x){\color{white}{\LARGE|}}\right] ^2} }\)

OpenStudy (solomonzelman):

for example, when you take a derivative of an absolute value, this is very helpful.